# Cut-and-paste for impulsive gravitational waves with $\Lambda$: The mathematical analysis

### v1

@@ 2 Nonexpanding impulsive waves with $\Lambda$
2.1 Metric representations & the ‘discontinuous transformation’
@@
where ${\mathcal{H}}$ is a real-valued function and $\delta$ is the Dirac-distribution. Due to the occurrence of a distributional coefficient, (2.1) lies far beyond the Geroch-Traschen class of metrics [12], which is defined by possessing Sobolev regularity $W^{1,2}_{\mbox{\small loc}}\cap L^{\infty}_{\mbox{\small loc}}$. It is known to be the largest class which allows in general to stably define the curvature in distributions (see also [27, 50]). Nevertheless, due to its simple structure the curvature of (2.1) can be computed explicitly to give the impulsive Newman-Penrose components $\Psi_{4}=(1+\frac{1}{6}\Lambda\eta\bar{\eta})^{2}{\mathcal{H}}_{,\eta\eta}\delta({{\mathcal{U}}})$, and $\Phi_{22}=(1+\frac{1}{6}\Lambda\eta\bar{\eta})\big{(}(1+\frac{1}{6}\Lambda\eta\bar{\eta}){\mathcal{H}}_{,\eta\bar{\eta}}+\frac{1}{6}\Lambda({\mathcal{H}}-\eta{\mathcal{H}}_{,\eta}-\bar{\eta}{\mathcal{H}}_{,\bar{\eta}})\big{)}\delta({{\mathcal{U}}})$.
@@ 2.2 Results in the pp-wave case
A first rigorous result in this realm has been established in [21] in the special case of impulsive pp-waves. There nonlinear distributional geometry [16, Chapter 3] based on algebras of generalized functions [7] has been employed to show the following: The ‘discontinuous coordinate change’ (2.5) relating the distributional Brinkmann form of the metric, i.e., (2.1) with $\Lambda=0$ to the continuous Rosen form, i.e., (2.2) with $\Lambda=0$ is the distributional limit of a generalized diffeomorphism, a concept to be detailed below. Intuitively speaking this approach consists in viewing the impulsive wave as a limiting case of a sandwich wave with an arbitrarily regularized wave profile, where the two forms of the metric arise as its (distributional) limits taken in different coordinate systems. This result rests on two pillars:
@@ 2.3 The five-dimensional formalism
and the associated real coordinates $x,y$ with $\eta=1/\sqrt{2}(x+iy)$. Finally the profile functions are related by
@@ 3.1 Nonlinear distributional geometry
The theory we are going to summarize (for all details see [22, 23], [16, Section 3.2]) rests on J.F. Colombeau’s construction of (so-called special) algebras of generalized functions [7]. These provide an extension of the linear theory of Schwartz distributions to the nonlinear realm retaining maximal consistency with classical analysis. The basic idea of the construction is regularization of distributions via nets of smooth functions combined with asymptotic estimates in terms of a regularization parameter.
On a smooth (second countable and Hausdorff) manifold $M$ denote by ${\mathcal{E}}(M)$ the set of all nets of smooth functions $(u_{\varepsilon})_{\varepsilon\in(0,1]=:I}$ which in addition depend smoothly [^3] Smooth dependence on the parameter renders the theory technically more pleasant but was not assumed in earlier references, for details see [5, Section 1]. on $\varepsilon$. The algebra of generalized functions on $M$ is defined as the quotient ${\mathcal{G}}(M):={\mathcal{E}}_{\mathrm{M}}(M)/{\mathcal{N}}(M)$ of moderate modulo negligible nets in ${\mathcal{E}}(M)$, which are defined via the following asymptotic estimates
@@
Here ${\mathcal{P}}$ denotes the space of all linear differential operators on $M$ and $K\Subset M$ means that $K$ is a compact subset of $M$. We write $u=[(u_{\varepsilon})_{\varepsilon}]$ for the elements of ${\mathcal{G}}(M)$ and call $(u_{\varepsilon})_{\varepsilon}$ a representative of the generalized function $u$. With sums, products, and the Lie derivative defined componentwise (i.e., for fixed $\varepsilon$) ${\mathcal{G}}(M)$ becomes a fine sheaf of differential algebras.
@@
The space of distributions $\mathcal{D}^{\prime}(M)$ can be embedded into ${\mathcal{G}}(M)$ via sheaf homomorphisms $\iota$ that preserve the product of $\mathcal{C}^{\infty}(M)$-functions. A coarser way of relating generalized functions in ${\mathcal{G}}(M)$ to distributions is as follows: $u\in{\mathcal{G}}(M)$ is called associated with $v\in{\mathcal{G}}(M)$, written $u\approx v$, if $u_{\varepsilon}-v_{\varepsilon}\to 0$ in $\mathcal{D}^{\prime}(M)$. Moreover, $w\in\mathcal{D}^{\prime}(M)$ is called associated with $u$ if $u\approx\iota(w)$.
@@
More generally the space of generalized sections of a vector bundle $E\to M$ is defined as $\Gamma_{\mathcal{G}}(M,E)={\mathcal{G}}(M)\otimes_{\mathcal{C}^{\infty}(M)}\Gamma(M,E)=L_{\mathcal{C}^{\infty}(M)}(\Gamma(M,E^{*}),{\mathcal{G}}(M)).$ It is a fine sheaf of finitely generated and projective ${\mathcal{G}}$-modules. For generalized tensor fields of rank $r,s$ we use the notation
${\mathcal{G}}^{r}_{s}(M)\cong L_{{\mathcal{G}}(M)}({\mathcal{G}}^{1}_{0}(M)^{s},{\mathcal{G}}^{0}_{1}(M)^{r};{\mathcal{G}}(M)).$ (3.1)
@@
Observe that it is possible to insert generalized vector fields and one-forms into generalized tensors, which is not possible in the distributional setting, cf. [28, 8]. This in turn allows to work with generalized metrics much as in the smooth setting. Here a generalized pseudo-Riemannian metric is a section $g\in{\mathcal{G}}^{0}_{2}(M)$ that is symmetric with determinant $\det g$ invertible in ${\mathcal{G}}$ (equivalently $|\det(g_{\varepsilon})_{ij}|\geq\varepsilon^{m}$ for some $m$ on compact sets), and a well-defined index $\nu$ (the index of $g_{\varepsilon}$ equals $\nu$ for $\varepsilon$ small). By a ‘globalization Lemma’ in [25, Lemma 2.4, p. 6] any generalized metric $g$ possesses a representative $(g_{\varepsilon})_{\varepsilon}$ such that each $g_{\varepsilon}$ is a smooth metric globally on $M$. We call a pair $(M,g)$ consisting of a smooth manifold and a generalized pseudo-Riemannian (Lorentzian) metric a generalized pseudo-Riemannian (Lorentzian) manifold, and a generalized spacetime if, in addition to being Lorentzian, it can be time oriented by a smooth vector field. This setting consistently extends the ‘maximal distributional’ one of Geroch and Traschen, see [50, 48]. In particular, any generalized metric induces an isomorphism between generalized vector fields and one-forms, and there is a unique Levi-Civita connection $\nabla$ corresponding to $g$.
@@
Next, to speak of geodesics one uses the space of generalized curves ${\mathcal{G}}[J,M]$ taking values in $M$, defined on an interval $J$. It is again a quotient of moderate modulo negligible nets $(\gamma_{\varepsilon})_{\varepsilon}$ of smooth curves, where we call a net moderate (negligible) if $(\psi\circ\gamma_{\varepsilon})_{\varepsilon}$ is moderate (negligible) for all smooth $\psi:M\to\mathbb{R}$. In addition, $(\gamma_{\varepsilon})_{\varepsilon}$ is supposed to be c-bounded, which means that $\gamma_{\varepsilon}(K)$ is contained in a compact subset of $M$ for $\varepsilon$ small and all compact sets $K\Subset J$. Observe that no distributional counterpart of such a space exists and it has long been realized that regularization is a possible remedy, cf. [28].
@@
The induced covariant derivative of a generalized vector field $\xi=[(\xi_{\varepsilon})_{\varepsilon}]\in{\mathcal{G}}^{1}_{0}(M)$ on a generalized curve $\gamma=[(\gamma_{\varepsilon})_{\varepsilon}]\in{\mathcal{G}}[J,M]$ is defined componentwise (i.e., by the classical fomulae for fixed $\varepsilon$) and gives again a generalized vector field $\xi^{\prime}$ on $\gamma$. In particular, a geodesic in a generalized spacetime is a curve $\gamma\in{\mathcal{G}}[J,M]$ satisfying $\gamma^{\prime\prime}=0$. Equivalently the usual local formula holds, i.e.,
@@
where $\Gamma^{k}_{ij}=[({\Gamma_{\varepsilon}}^{k}_{ij})_{\varepsilon}]$ denotes the Christoffel symbols of the generalized metric $g=[(g_{\varepsilon})_{\varepsilon}]$. Finally we say that a generalized spacetime $(M,g)$ is geodesically complete if every geodesic $\gamma$ can be defined on $\mathbb{R}$ [42, Definition 2.1, p. 240].
3.2 Impulsive waves as generalized spacetimes & the geodesic equation
In this section we introduce the generalized metric form of nonexpanding impulsive waves for arbitrary values of $\Lambda$, using the five-dimensional formalism of Section 2.3. Indeed, starting with the metric (2.6) we replace the Dirac-delta with a generic regularization: Choose any smooth function $\rho$ on $\mathbb{R}$ with unit integral and support in $[-1,1]$ and for $\varepsilon\in(0,1]$ set $\delta_{\varepsilon}(x):=(1/\varepsilon)\,\rho(x/\varepsilon)$. Such a net $(\delta_{\varepsilon})_{\varepsilon}$ is called a model delta net and we use it to define the regularized pseudo-Riemannian manifold $(\bar{M}=\mathbb{R}^{5},\bar{g}_{\varepsilon})$ with line element
@@ 3.2 Impulsive waves as generalized spacetimes & the geodesic equation
Hence $(\bar{M},\bar{g}_{\varepsilon})$ is a smooth sandwich wave which is flat space outside the wave zone given by $|U|\leq\varepsilon$. The regularized impulsive wave spacetime of our interest $(M,g_{\varepsilon})$ is now given by the (anti-)de Sitter hyperboloid (2.7) embedded in $(\bar{M},\bar{g}_{\varepsilon})$.
To obtain an impulsive wave metric in ${\mathcal{G}}^{0}_{2}(\mathbb{R}^{5})$ we use a model delta function, that is an element $D\in{\mathcal{G}}(\mathbb{R})$ that has a model delta net as a representative, $D=[(\delta_{\varepsilon})_{\varepsilon}]$. Next we consider the $5$-dimensional generalized impulsive pp-wave manifold $(\bar{M}=\mathbb{R}^{5},\bar{g})$ with
@@
One easily checks that this defines a generalized metric with representative (3.3). At this point we specify the (A)dS hyperboloid $M$ in $(\bar{M},\bar{g})$ as usual, explcitly by
@@
Note that $M$ is a (classical) smooth hypersurface. Finally, we restrict the metric $\bar{g}$ (again componentwise, that is for fixed $\varepsilon$) to $M$ to obtain the generalized spacetime $(M,g)$ which we take as our model of nonexpanding impulsive waves propagating in a(n anti-)de Sitter universe.
To derive the geodesic equations in $(M,g$) we use the fact that in nonlinear generalized functions all classical formulae hold for fixed $\varepsilon$. So we derive the $M$-geodesics from the condition that their $\bar{M}$-acceleration is normal to $M$, ¯$\bar{\nabla}_{T}T=-\sigma g(T,\bar{\nabla}_{T}N)N/g(N,N)$. Here $\bar{\nabla}$ is the generalized Levi-Civita connection of $(\bar{M},\bar{g})$, and $T$ and $N$ denote the geodesic tangent and the (non-normalized) normal vector to $M$ defined via its representative $N^{\alpha}_{\varepsilon}=g^{\alpha\beta}_{\varepsilon}dF_{\beta}$, respectively. In this way we arrive at the geodesic equations for $\gamma=(U,V,Z_{p})$:
@@ 3.3 Existence and uniqueness of geodesics
Next we briefly indicate how one proves unique solvability of the initial value problem for differential equations like (3.2) in generalized functions. This is basically done in three steps:
@@
    Observe that a solution candidate $(\gamma_{\varepsilon})_{\varepsilon}$ actually is comprised of geodesics $\gamma_{\varepsilon}$ of the regularized spacetime $(M,g_{\varepsilon})$, cf. (3.3).
@@
    One shows existence of a generalized solution by establishing c-boundedness and moderateness of the solution candidate, i.e., $\gamma:=[(\gamma_{\varepsilon})_{\varepsilon}]\in{\mathcal{G}}[J,M]$.
@@
We will refer to $\gamma^{-}$ as seed geodesics and start to think of it as geodesics in the impulsive wave spacetime (2.6), (2.7) ‘in front’ of the impulse that is for $U^{-}<0$. Also, $\gamma^{-}$ is a geodesic in the regularized spacetime (3.3), (2.7) ‘in front’ of the sandwich wave, that is for $U^{-}\leq-\varepsilon$. We will denote the affine parameter time when $\gamma^{-}$ enters this regularization wave region by $\alpha_{\varepsilon}$, i.e.,
@@
Observe that $\alpha_{\varepsilon}\to 0$ from below as $\varepsilon\to 0$. Finally, we come to set up the data for the solution candidate $\gamma_{\varepsilon}$ of the system (3.2) by
@@
Figure 1: The $U$-component of the seed geodesic $\gamma^{-}$ is depicted in black until it reaches the regularisation sandwich at parameter time $\lambda=\alpha_{\varepsilon}$, i.e., $U^{-}(\alpha_{\varepsilon})=-\varepsilon$. While in the background spacetime it would continue as the dotted red line to $U=0$ at $\lambda=0$, in the regularised spacetime it continues as solution $\gamma_{\varepsilon}$ of ((1)) with data (3.13) (depicted in green). Theorem 3.1 guarantees that $\gamma_{\varepsilon}$ (for $\varepsilon$ small) leaves the regularisation sandwich at $\lambda=\beta_{\varepsilon}$ and continues as background geodesic.
The central result on the solvability of the geodesic equations for the generalized spacetime (3.4), (3.2) now is the following, cf. [43, Thms. 3.6, 3.7].

### v2

@@ 2 Nonexpanding impulsive waves with $\Lambda$
2.1 Metric representations and the ‘discontinuous transformation’
@@
where ${\mathcal{H}}$ is a real-valued function and $\delta$ is the Dirac-distribution. Due to the occurrence of a distributional coefficient, (2.1) lies far beyond the Geroch-Traschen class of metrics [12], which is defined by possessing Sobolev regularity $W^{1,2}_{\mbox{\small loc}}\cap L^{\infty}_{\mbox{\small loc}}$. It is known to be the largest class which allows one in general to stably define the curvature in distributions (see also [27, 50]). Nevertheless, due to its simple structure the curvature of (2.1) can be computed explicitly to give the impulsive Newman-Penrose components $\Psi_{4}=(1+\frac{1}{6}\Lambda\eta\bar{\eta})^{2}{\mathcal{H}}_{,\eta\eta}\delta({{\mathcal{U}}})$, and $\Phi_{22}=(1+\frac{1}{6}\Lambda\eta\bar{\eta})\big{(}(1+\frac{1}{6}\Lambda\eta\bar{\eta}){\mathcal{H}}_{,\eta\bar{\eta}}+\frac{1}{6}\Lambda({\mathcal{H}}-\eta{\mathcal{H}}_{,\eta}-\bar{\eta}{\mathcal{H}}_{,\bar{\eta}})\big{)}\delta({{\mathcal{U}}})$.
@@ 2.2 Results in the pp-wave case
A first rigorous result in this realm has been established in [21] in the special case of impulsive pp-waves. There, nonlinear distributional geometry [16, Chapter 3] based on algebras of generalised functions [7] has been employed to show the following: The ‘discontinuous coordinate change’ (2.5) relating the distributional Brinkmann form of the metric, i.e., (2.1) with $\Lambda=0$ to the continuous Rosen form, i.e., (2.2) with $\Lambda=0$ is the distributional limit of a generalised diffeomorphism, a concept to be detailed below. Intuitively speaking this approach consists in viewing the impulsive wave as a limiting case of a sandwich wave with an arbitrarily regularised wave profile, where the two forms of the metric arise as its (distributional) limits taken in different coordinate systems. This result rests on two pillars:
@@ 2.3 The five-dimensional formalism
and the associated real coordinates $x,y$ with $\eta=1/\sqrt{2}(x+iy)$. Finally, the profile functions are related by
@@ 3.1 Nonlinear distributional geometry
The theory we are going to summarize (for all details see [22, 23], [16, Section 3.2]) rests on J.F. Colombeau’s construction of (so-called special) algebras of generalised functions [7]. These provide an extension of the linear theory of Schwartz distributions to the nonlinear realm retaining maximal consistency with classical analysis. The basic idea of the construction is regularisation of distributions via nets of smooth functions combined with asymptotic estimates in terms of a regularisation parameter.
On a smooth (second countable and Hausdorff) manifold $M$ denote by ${\mathcal{E}}(M)$ the set of all nets of smooth functions $(u_{\varepsilon})_{\varepsilon\in(0,1]=:I}$ which in addition depend smoothly [^3] Smooth dependence on the parameter renders the theory technically more pleasant but was not assumed in earlier references, for details see [5, Section 1]. on $\varepsilon$. The algebra of generalised functions on $M$ is defined as the quotient ${\mathcal{G}}(M):={\mathcal{E}}_{\mathrm{M}}(M)/{\mathcal{N}}(M)$ of moderate modulo negligible nets in ${\mathcal{E}}(M)$, which are defined via the following asymptotic estimates
@@
Here ${\mathcal{P}}$ denotes the space of all linear differential operators on $M$ and $K\Subset M$ means that $K$ is a compact subset of $M$. We write $u=[(u_{\varepsilon})_{\varepsilon}]$ for the elements of ${\mathcal{G}}(M)$ and call $(u_{\varepsilon})_{\varepsilon}$ a representative of the generalised function $u$. With sums, products, and the Lie derivative defined componentwise (i.e., for fixed $\varepsilon$) ${\mathcal{G}}(M)$ becomes a fine sheaf of differential algebras.
@@
The space of distributions $\mathcal{D}^{\prime}(M)$ can be embedded into ${\mathcal{G}}(M)$ via sheaf homomorphisms $\iota$ that preserve the product of $\mathcal{C}^{\infty}(M)$-functions. A coarser way of relating generalised functions in ${\mathcal{G}}(M)$ to distributions is as follows: $u\in{\mathcal{G}}(M)$ is called associated with $v\in{\mathcal{G}}(M)$, written $u\approx v$, if $u_{\varepsilon}-v_{\varepsilon}\to 0$ in $\mathcal{D}^{\prime}(M)$. Moreover, $w\in\mathcal{D}^{\prime}(M)$ is called associated with $u$ if $u\approx\iota(w)$.
@@
More generally the space of generalised sections of a vector bundle $E\to M$ is defined as $\Gamma_{\mathcal{G}}(M,E)={\mathcal{G}}(M)\otimes_{\mathcal{C}^{\infty}(M)}\Gamma(M,E)=L_{\mathcal{C}^{\infty}(M)}(\Gamma(M,E^{*}),{\mathcal{G}}(M)).$ It is a fine sheaf of finitely generated and projective ${\mathcal{G}}$-modules. For generalised tensor fields of rank $r,s$ we use the notation
${\mathcal{G}}^{r}_{s}(M)\cong L_{{\mathcal{G}}(M)}({\mathcal{G}}^{0}_{1}(M)^{r},{\mathcal{G}}^{1}_{0}(M)^{s};{\mathcal{G}}(M)).$ (3.1)
@@
Observe that it is possible to insert generalised vector fields and one-forms into generalised tensors, which is not possible in the distributional setting, cf. [28, 8]. This in turn allows one to work with generalised metrics much as in the smooth setting. Here a generalised pseudo-Riemannian metric is a section $g\in{\mathcal{G}}^{0}_{2}(M)$ that is symmetric with determinant $\det g$ invertible in ${\mathcal{G}}$ (equivalently $|\det(g_{\varepsilon})_{ij}|\geq\varepsilon^{m}$ for some $m$ on compact sets), and a well-defined index $\nu$ (the index of $g_{\varepsilon}$ equals $\nu$ for $\varepsilon$ small). By a ‘globalization Lemma’ in [25, Lemma 2.4, p. 6] any generalised metric $g$ possesses a representative $(g_{\varepsilon})_{\varepsilon}$ such that each $g_{\varepsilon}$ is a smooth metric globally on $M$. We call a pair $(M,g)$ consisting of a smooth manifold and a generalised pseudo-Riemannian (Lorentzian) metric a generalised pseudo-Riemannian (Lorentzian) manifold, and a generalised spacetime if, in addition to being Lorentzian, it can be time oriented by a smooth vector field. This setting consistently extends the ‘maximal distributional’ one of Geroch and Traschen, see [50, 48]. In particular, any generalised metric induces an isomorphism between generalised vector fields and one-forms, and there is a unique Levi-Civita connection $\nabla$ corresponding to $g$.
@@
Next, to speak of geodesics one uses the space of generalised curves ${\mathcal{G}}[J,M]$ taking values in $M$, defined on an interval $J$. It is again a quotient of moderate modulo negligible nets $(\gamma_{\varepsilon})_{\varepsilon}$ of smooth curves, where we call a net moderate (negligible) if $(\psi\circ\gamma_{\varepsilon})_{\varepsilon}$ is moderate (negligible) for all smooth $\psi:M\to\mathbb{R}$. In addition, $(\gamma_{\varepsilon})_{\varepsilon}$ is supposed to be c-bounded, which means that $\gamma_{\varepsilon}(K)$ is contained in a compact subset of $M$ for $\varepsilon$ small and all compact sets $K\Subset J$. Observe that no distributional counterpart of such a space exists and it has long been realised that regularisation is a possible remedy, cf. [28].
@@
The induced covariant derivative of a generalised vector field $\xi=[(\xi_{\varepsilon})_{\varepsilon}]\in{\mathcal{G}}^{1}_{0}(M)$ on a generalised curve $\gamma=[(\gamma_{\varepsilon})_{\varepsilon}]\in{\mathcal{G}}[J,M]$ is defined componentwise (i.e., by the classical fomulae for fixed $\varepsilon$) and gives again a generalised vector field $\xi^{\prime}$ on $\gamma$. In particular, a geodesic in a generalised spacetime is a curve $\gamma\in{\mathcal{G}}[J,M]$ satisfying $\gamma^{\prime\prime}=0$. Equivalently the usual local formula holds, i.e.,
@@
where $\Gamma^{k}_{ij}=[({\Gamma_{\varepsilon}}^{k}_{ij})_{\varepsilon}]$ denotes the Christoffel symbols of the generalised metric $g=[(g_{\varepsilon})_{\varepsilon}]$. Finally we say that a generalised spacetime $(M,g)$ is geodesically complete if every geodesic $\gamma$ can be defined on $\mathbb{R}$ [42, Definition 2.1, p. 240].
3.2 Impulsive waves as generalised spacetimes & the geodesic equation
In this section we introduce the generalised metric form of nonexpanding impulsive waves for arbitrary values of $\Lambda$, using the five-dimensional formalism of Section 2.3. Indeed, starting with the metric (2.6) we replace the Dirac-delta with a generic regularisation: Choose any smooth function $\rho$ on $\mathbb{R}$ with unit integral and support in $[-1,1]$ and for $\varepsilon\in(0,1]$ set $\delta_{\varepsilon}(x):=(1/\varepsilon)\,\rho(x/\varepsilon)$. Such a net $(\delta_{\varepsilon})_{\varepsilon}$ is called a model delta net and we use it to define the regularised pseudo-Riemannian manifold $(\bar{M}=\mathbb{R}^{5},\bar{g}_{\varepsilon})$ with line element
@@ 3.2 Impulsive waves as generalised spacetimes & the geodesic equation
Hence $(\bar{M},\bar{g}_{\varepsilon})$ is a smooth sandwich wave which is flat space outside the wave zone given by $|U|\leq\varepsilon$. The regularised impulsive wave spacetime of our interest $(M,g_{\varepsilon})$ is now given by the (anti-)de Sitter hyperboloid (2.7) embedded in $(\bar{M},\bar{g}_{\varepsilon})$.
To obtain an impulsive wave metric in ${\mathcal{G}}^{0}_{2}(\mathbb{R}^{5})$ we use a model delta function, that is an element $D\in{\mathcal{G}}(\mathbb{R})$ that has a model delta net as a representative, $D=[(\delta_{\varepsilon})_{\varepsilon}]$. Next we consider the $5$-dimensional generalised impulsive pp-wave manifold $(\bar{M}=\mathbb{R}^{5},\bar{g})$ with
@@
One easily checks that this defines a generalised metric with representative (3.3). At this point we specify the (A)dS hyperboloid $M$ in $(\bar{M},\bar{g})$ as usual, explcitly by
@@
Note that $M$ is a (classical) smooth hypersurface. Finally, we restrict the metric $\bar{g}$ (again componentwise, that is for fixed $\varepsilon$) to $M$ to obtain the generalised spacetime $(M,g)$ which we take as our model of nonexpanding impulsive waves propagating in a(n anti-)de Sitter universe.
To derive the geodesic equations in $(M,g$) we use the fact that in nonlinear generalised functions all classical formulae hold for fixed $\varepsilon$. So we derive the $M$-geodesics from the condition that their $\bar{M}$-acceleration is normal to $M$, ¯$\bar{\nabla}_{T}T=-\sigma g(T,\bar{\nabla}_{T}N)N/g(N,N)$. Here $\bar{\nabla}$ is the generalised Levi-Civita connection of $(\bar{M},\bar{g})$, and $T$ and $N$ denote the geodesic tangent and the (non-normalised) normal vector to $M$ defined via its representative $N^{\alpha}_{\varepsilon}=g^{\alpha\beta}_{\varepsilon}dF_{\beta}$, respectively. In this way we arrive at the geodesic equations for $\gamma=(U,V,Z_{p})$:
@@ 3.3 Existence and uniqueness of geodesics
Next we briefly indicate how one proves unique solvability of the initial value problem for differential equations like (3.2) in generalised functions. This is basically done in three steps:
@@
    Observe that a solution candidate $(\gamma_{\varepsilon})_{\varepsilon}$ actually is comprised of geodesics $\gamma_{\varepsilon}$ of the regularised spacetime $(M,g_{\varepsilon})$, cf. (3.3).
@@
    One shows existence of a generalised solution by establishing c-boundedness and moderateness of the solution candidate, i.e., $\gamma:=[(\gamma_{\varepsilon})_{\varepsilon}]\in{\mathcal{G}}[J,M]$.
@@
We will refer to $\gamma^{-}$ as seed geodesics and start to think of it as geodesics in the impulsive wave spacetime (2.6), (2.7) ‘in front’ of the impulse, that is for $U^{-}<0$. Also, $\gamma^{-}$ is a geodesic in the regularised spacetime (3.3), (2.7) ‘in front’ of the sandwich wave, that is for $U^{-}\leq-\varepsilon$. We will denote the affine parameter time when $\gamma^{-}$ enters this regularisation wave region by $\alpha_{\varepsilon}$, i.e.,
@@
Observe that $\alpha_{\varepsilon}\to 0$ from below as $\varepsilon\to 0$. Finally, we come to setting up the data for the solution candidate $\gamma_{\varepsilon}$ of the system (3.2) by
@@
$\gamma^{-}=\gamma_{\varepsilon}$
$\gamma^{-}$
$\gamma_{\varepsilon}$
$\gamma^{+}_{\varepsilon}=\gamma_{\varepsilon}$
00
Figure 1: The $U$-component of the seed geodesic $\gamma^{-}$ is depicted in black until it reaches the regularisation sandwich at parameter time $\lambda=\alpha_{\varepsilon}$, i.e., $U^{-}(\alpha_{\varepsilon})=-\varepsilon$. While in the background spacetime it would continue as the dotted red line to $U=0$ at $\lambda=0$, in the regularised spacetime it continues as a solution $\gamma_{\varepsilon}$ of ((1)) with data (3.13) (depicted in green). Theorem 3.1 guarantees that $\gamma_{\varepsilon}$ (for $\varepsilon$ small) leaves the regularisation sandwich at $\lambda=\beta_{\varepsilon}$ and continues as a background geodesic.
The central result on the solvability of the geodesic equations for the generalised spacetime (3.4), (3.2) now is the following, cf. [43, Thms. 3.6, 3.7].